[Bteam] notes from last Wednesday's meeting

Arne Freyberger freyberg at jlab.org
Mon Dec 13 15:41:58 EST 2010


Also attached are my running notes on the Arc1 optics, see final entry 
of these notes for some late breaking news.  The prioritized BSY 
diagnostics package for FY11 is also attached.
Arne


  B-team Meeting Notes


    1 2010-12-08 Wed 14:00 BTeam Meeting--Notes


      1.1 Optics-on-Call Report Kazimi Hofler

No calls received. Support Injector transport after the cessiate-bake 
effort on Tuesday. 150? A setup complete with more interception than 
previous setup.


      1.2 Beam Studies Report

    * Summary of Heat-Activate
      Pockel Cell alignment did not take place. Plenty of plenary steps,
      but no alignment.

      Re-thinking the requirements for this phase of the QWeak
      experiment. Would like a few days of the Winter break. EGG group
      will support this. What else do we need?

      Next heat-cessiate party will be during the Winter Break.

    * Parity Quality Beam Status Report Qweak
      Charge asymmetry data. Short periods where the A_Q is not
      controlled due to data acquisition issues. But when the feedback
      loop is working A_Q is reasonably stable.

      Position differences are small when the A_Q is stable. Future
      position differences will be for 1h slugs of data. Expect the
      differences to be in the 10s of nm. Present position differences
      good enough for this phase of the experiment.

      Present asymmetries look encouraging for the run out through May
      13^th

          o Compton Signal
            Signal to background looks very good except that there is a
            20minute rise and fall of the background. Not sure where the
            background comes from.

            Speculation is that it is correlated with the QE hole on the
            cathode.

          o Things that still need to be checked out
                + modulation system and hall optics properly configure
                + BCM noise issue (EES involved)
                + Measure betatron damping via PZT
                + optimize FFB FF system, what are the requirements? If
                  the asymmetries look ok do we really need to improve
                  the FFB FF system? "I would not mess with it", says
                  Roger Carlini.
                + IA feedback
                + 0.2 Hz kick? Still present in the Injector. Hall needs
                  to evaluate the impact of this and report back. This
                  has been a nagging issue since PRex.


      1.3 Beam Studies Planning

    * RF Recovery
    * High Current (180 {?}A) Injector setup for QWeak
    * Mott Measurements
    * Arc9 Transport


      1.4 BSY Diagnostics Benesch

Present updated table. B-team to prioritize the list based on 
desirability and known resource limits. The outcome of this discussion 
will be the plan that is executed.

See Jay's attached list with the prioritized list.

    * TODO Need to finish ItoV on the dumplettes


      1.5 12GeV Pathlength Requirements Status Tiefenback

Update on requirements table and overall documentation preparation.


      1.6 ARC1 model Arne

Results of the latest round of Arc1 modeling. See separate note. Recent 
developments have identified the problem with getting the Arc1 transport 
to be symmetric. The body gradient for MBE1A01 is unique when compared 
to the other 15dipoles in the arc which all have the same body gradient. 
This problem is in the Art++ decks as well. See separate note for details.

The 12 GeV decks have used the MBE1A01 value for all 16 dipoles in the 
Arc, whereas a better choice might be to set MBE1A01 to the values found 
for the other 15 dipoles. Waiting guidance on this....


    2 2010-12-01 Wed 14:00 BTeam Meeting--Notes


      2.1 Optics-on-Call Report: Hofler

Summary of the thankful week

Fri: Hall C Compton scan -- done.

Weekend: Energy stability problems, gang phase lock still a problem? A 
power cycle may be required; readback may be invalid; also an observed 
current dependence in tracking.

Wed: Hall C Compton backed out, halo counts up, fixed by rephasing. Need 
to get back through Compton. Still an open issue, should revisit after 
spot move on Thursday 12-02.


      2.2 Beam Studies Report: All

    * Pathlength Analysis: Yunn
      Status of the analysis effort of the tunnel heating/pathlength
      five-pass test in September, final report due next week. Byung's
      draft report is attached. Fig 1 shows slower cooling than warming.
      12 GeV conditions will likely give RF shift of 2-3 degrees/arc.
      15F LCW rise was half of 12 GeV conditions; concrete temp did not
      asymptote. Should we reperform this study with a longer test
      and/or higher temperature? Is this enough data to draw conclusions?

      Should we get a baseline to separate out baseline pathlength
      changes vs temperature-driven changes? This is small compared to
      expected effect. Estimate that these are acceptable but more
      numbers on air/wall temp and air conditioning required. Should
      pull sensors for data and compare, and plan to deliver conclusions
      before winter break (Tief).

    * Arc9 model calibration test: Roblin
      ATLiS entry 10511
      <http://opsweb.acc.jlab.org/CSUEApps/atlis/task/10511>. Multiple
      RF FSD trips, ran fopt three times, not necessarily complete,
      requires data quality review by Yves.
    * Hall C double peak source found/resolved? Suleiman
      See this elog entry
      <http://opweb.acc.jlab.org/CSUEApps/elog/entry/1555706?PHPSESSID=eba0eb97d22d5f6c7ee0b57e9c387545>.
      Can we get a troubleshooting/diagnostic? Or have we lost some
      troubleshooting for low current by turning off the beam scraping
      monitor? It is a useful archived diagnostic and should evaluate
      whether we can turn it back on in future. We should ensure that it
      stays off until we want it back on. Was also used as low-level
      transfer function monitor, monitoring for changes in lattice. We
      will return to this in the future to see if it can be acceptably
      turned back on. We'll modify app and send signals to Qweak for
      monitoring in the meantime.


      2.3 Beam Studies Planning: All

    * Thursday 2010-12-02 and Tuesday 2010-12-07 Plans: Suggested
          o Tuesday 2010-12-07
            CHL power switch uncertainties, injector cathode activation
            (~12h). TPE also scheduled to start operations at start of
            evening shift.
          o Compton backgrounds (hall A)
          o Injector development for Hall C/Qweak (Dalton, Benesch), ~4h
            (to 12h?)
            ATLiS entry 10271
            <http://opsweb.acc.jlab.org/CSUEApps/atlis/task/10271>,
            Rebalance systematics, RHWP scans/PC translation scans. Was
            done before but with IHWP in wrong orientation. Oops! May
            require some time both days.

            Estimate 4h Thu 12-02, 8h Tue 12-07 (laser table in
            conjunction with Cathode activation, total time 12h)

          o One pass stripline BPM test, ~1h
            ATLiS entry 10445
            <http://opsweb.acc.jlab.org/CSUEApps/atlis/task/10445>,
            Large dynamic range of currents (100uA to >=100 nA CW) for
            BPM linearity test. Access required, not time critical; can
            wait for next access.
          o Repeat of Arc9 model calibration test?, ~2h
            ATLiS entry 10511,
            <http://opsweb.acc.jlab.org/CSUEApps/atlis/task/10511>
            should wait for Yves evaluation of existing data to see if
            its enough for his analysis.
          o Cryomodule JT valve test?, ~10h
            ATLiS entry 7874
            <http://opsweb.acc.jlab.org/CSUEApps/atlis/task/7874>, Has
            been lingering since June. Requires beam off. Also requires
            Jonathan Creel's availability.
    * Thursday 2010-12-02 Plans: Decided
          o Move spot (0.5h)
          o Injector development for QWeak (4h)
          o Compton background tuning (0.5-1h)
    * Tuesday 2010-12-07 Plans: Decided
          o Injector development for QWeak, laser table, cathode
            activation (8-12h)


      2.4 FY11 BTeam Projects

    * BSY Diagnostics: Benesch
      Details of the FY11 BSY diagnostics effort: showed list of 5
      diagnostics suggestions on overhead (FY11 diagnostic suggestions).
      Discussions of whether to relocate harps back to E03 crosses, or
      install new crosses in front of S10 girders for optics matching
      after each recombiner for 12 GeV upgrade commissioning development.
          o TODO ATLiS to test optics measurements at 1E03/2S10 (before
            Jan 2011 optics)
          o TODO Work up detailed tabular strawman, cost/resource
            estimates (Benesch)
          o TODO Cross-check injector harp optics match measurement
            consistency


      2.5 Anything else?

    * Optics: Hofler, Kazimi

Date: 2010-12-13 15:37:41 EST

HTML generated by org-mode 6.33x in emacs 23


-- 
Director of Accelerator Operations
Jefferson Lab           Office: (757) 269-6268
12000 Jefferson Avenue  Cell:   (757) 876-6289
Newport News, VA 23606  Email:  freyberg at jlab.org

-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/bteam/attachments/20101213/10708be7/attachment-0004.html>
-------------- next part --------------
A non-text attachment was scrubbed...
Name: FY11_diagnostic_suggestions.doc
Type: application/msword
Size: 43008 bytes
Desc: not available
URL: <https://mailman.jlab.org/pipermail/bteam/attachments/20101213/10708be7/attachment-0002.doc>
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/bteam/attachments/20101213/10708be7/attachment-0005.html>
-------------- next part --------------
A non-text attachment was scrubbed...
Name: ARC1.pdf
Type: application/pdf
Size: 1708793 bytes
Desc: not available
URL: <https://mailman.jlab.org/pipermail/bteam/attachments/20101213/10708be7/attachment-0002.pdf>


More information about the BTeam mailing list